jigglypuff314 (jigglypuff314):

dy/dx = tany / x I need to find the anti-derivative of that, pwease help :3

OpenStudy (abb0t):

use separation of variables.

OpenStudy (anonymous):

separate the variables and integrate

jigglypuff314 (jigglypuff314):

mmm how would I do that? :/ like dy/dx = (tany)(1/x) ?

OpenStudy (anonymous):

same process used in the previous problem

jigglypuff314 (jigglypuff314):

really? but how would I do dy/tany = dx/x ? :/

OpenStudy (abb0t):

\(\sf \frac{dy}{dx}=\frac{tan(y)}{x} \Rightarrow \frac{1}{x}dx=\frac{1}{tan(y)}dx\)

OpenStudy (abb0t):

sorry, the [(tan(y)]\(^{-1}\) should be dy, not dx

OpenStudy (anonymous):

1/tany = coty dx/x = (1/x)dx, so what is the integral of cot(y) and what is the integral of 1/x?

OpenStudy (abb0t):

yes.

jigglypuff314 (jigglypuff314):

mmmm -1 / (1+ y^2) = lnx ? :3

OpenStudy (abb0t):

what?

jigglypuff314 (jigglypuff314):

the anti derivative? coty dy = -1 / (1+y^2) right?

jigglypuff314 (jigglypuff314):

no wait, nvm that would be arccot >,<

OpenStudy (abb0t):

No.

jigglypuff314 (jigglypuff314):

lol xD ignore what I was saying before :P ln|siny| = lnx + C ?

OpenStudy (anonymous):

right

jigglypuff314 (jigglypuff314):

soooo siny = e^(lnx + C) -> siny = C x ?

OpenStudy (anonymous):

right

jigglypuff314 (jigglypuff314):

then y = arcsin(Cx) ??? :/

OpenStudy (anonymous):

right

jigglypuff314 (jigglypuff314):

That's all? Thank you ^_^

OpenStudy (anonymous):

\[\int\limits \cot y dy=\int\limits \frac{ \cos ~y }{ \sin ~y }dy=\ln \left| \sin y \right|\]